### v1

@@ 4 Related Work
GNS applies the message passing architecture of graph convolution networks (GCN) to physical system simulation (Sanchez-Gonzalez et al., 2018, 2020; Pfaff et al., 2020; Wu et al., 2022; Allen et al., 2023). Poli et al. (2019) connect GNS with blackbox neural ODEs (GNODE) (see also Jin et al., 2022; Bishnoi et al., 2022; Wu et al., 2022; Allen et al., 2023; Li et al., 2022). MNODE focuses on causal relationships between features across time, rather than between spatial interactions; as such, we use directed rather than undirected graphs.
@@ 5.1 Motivation: Safely Managing Exercise in T1D
Type 1 Diabetes (T1D) is an autoimmune condition characterized by the destruction of insulin producing beta-cells. To manage glucose concentrations and prevent diabetes-related complications, intensive insulin therapy through externally delivered sources (e.g., insulin pumps) is required. Regular physical activity and exercise lead to numerous health benefits such as increased insulin sensitivity, weight management, and improved psychosocial well-being. However, for individuals with T1D, due to the inability to rapidly decrease circulating insulin concentrations, exercise can also increase the risk of hypoglycemia. Despite technological advancements such as continuous glucose monitoring (CGM) and automated insulin delivery (AID) systems, individuals with T1D still face significant challenges with managing glucose concentrations around exercise. Many adults with T1D are not meeting current exercise recommendations of at least 30-minutes of moderate-to-vigorous physical activity (MVPA) per day (Riddell et al., 2017), with fear of exercise-related hypoglycemia a leading barrier (Brazeau et al., 2008). To address these challenges, there is a pressing demand for precise, reliable models that can predict individualized glycemic responses during exercise and encourage safe physical activity for all individuals with T1D.
Past efforts on modeling glycemic response to physical activity have focused on developing more intricate mechanistic models (Dalla Man et al., 2009; Liu et al., 2018; Deichmann et al., 2023), rather than devising a performant predictive model. Although glucose prediction using ML and, more recently, deep learning methods has received significant attention, a dearth of papers consider exercise periods of exercise (Oviedo et al., 2017) and the few that do do not perform well (Hobbs et al., 2019; Xie and Wang, 2020; Tyler et al., 2022). We aim to leverage hybrid modeling to predict glucose concentrations of an individual with T1D during the first 30 minutes of a planned physical activity, given historical context and expected future covariates. We chose a 30-minute prediction window because most recorded exercises in our dataset (see Sec. 5.2) last for around 30 minutes. Modeling exercise-period glucose levels is already a challenging problem and we want to distinguish this task from the (even harder) task of the post-exercise period.
@@ 5.2 Data Preparation
Our data come from the Type 1 Diabetes Exercise Initiative (T1DEXI) (Riddell et al., 2023), which can be requested via https://doi.org/10.25934/PR00008428. This is a real-world study of exercise effects on 497 adults with T1D. Participants in the study were randomly assigned to one of three types of study-specific exercise videos: aerobic, resistance, or interval for a total of six sessions over four weeks. Participants were asked to self-report their food intake and exercise habits, while their insulin dosage and relevant physiological responses were recorded by corresponding wearable devices such as insulin pumps, CGM devices, and smart watches.
@@
For selected participants, we first filter out exercises shorter than 30 minutes. Then, for each exercise instance, we extract the participant’s metabolic history (basal/bolus insulin delivery, carbohydrate intake, heart rate, step count, and CGM readings) 4 hours before the start of exercise to 30 minutes after to form a 5-dimensional time series. The 4-hour context window was chosen because the effect of most bolus insulin (fast-acting insulin) lasts for around 4 hours. It is common within the T1D community to consider this time range when historical context is relevant in decision making. Finally, we filter out exercise instances with missing values. We end up with 94 exercise instances from 32 participants. See the Appendix for further data preprocessing details.
Counterfactual Simulation Sets
For each exercise instance, we form three copies to modify with one of the following sets of interventions, selected uniformly at random: (1) adding 0/50/100 grams (g) of carbohydrate at the beginning of exercise; (2) adding 0/2.5/5.0 units of insulin uniformly throughout the first 30 minutes of exercise; (3) no-change/50g carbs/10.0 units of insulin at the beginning of exercise; (4) replacing the real exercise heart rate trajectory with a prototypical trajectory seen in aerobic/resistance/interval training. These intervention sets capture the most integral pieces of well-understood and well-researched domain knowledge, allowing us to compute a reliable true ranking. In particular, increasing levels of carbohydrates progressively increase mean glucose, increasing levels of insulin progressively decrease mean glucose, and increasing exercise intensity (as measured by heart rate) leads to progressively increasing mean glucose during the short exercise period being considered (Riddell et al., 2017).
@@ 5.3 Key Implementation Details
and then computes the gradient via backpropagation.
In the above equation we use subscript instead of parenthesis to indicate the transition from the continuous time domain to discrete time steps. In our implementation, we set $\Delta t$ to be 5 minutes, which is sampling rate of CGM readings in the T1DEXI data.
@@
Instead of the full UVA/Padova S2013 model (Man et al., 2014), which suffers from over-parameterization and redundancy, our hybrid models build upon a reduced model with the aim of reducing variance and improving generalization performance. We applied a data-driven reduction method to reduce the full UVA/Padova mechanistic model and its causal graph. See the Appendix.
@@
We consider 3 hybrid models in the form of Eq. (5) with increasing amounts of flexibility: (1) latent parameter dynamics (LP) defined via $c_{1}=c_{4}=1,c_{2}=c_{3}=0$; (2) latent parameter dynamics plus state closure (LPSC) defined via $c_{1}=c_{2}=c_{4}=1,c_{3}=0$; and (3) mechanistic neural ODE (MNODE) defined via $c_{2}=1,c_{1}=c_{3}=c_{4}=0$. In all of these cases, the adjacency matrices $A_{s}$ and $A_{x}$ are specified based on our reduced UVA/Padova mechanistic model. We also consider the blackbox neural ODE model (BNODE) defined as in MNODE via $c_{2}=1,c_{1}=c_{3}=c_{4}=0$, but with full adjacency matrices $A_{s}=A_{x}=\mathbf{1}$ to allow arbitrary causal relationships to be learned. Finally, we provide a set of non-ODE-based blackbox sequence models: the LSTM, Transformer (Trans), temporal convolutional network (TCN), and diagonal approximation to S4 (S4D). For our mechanistic model baseline, we consider the full (not reduced) UVA-Padova S2013 simulator (UVA). We limit all models to have fewer than 25,000 parameters for fair comparison. For all of our ODE-based models, we use a multi-stack LSTM for our initial-condition encoder; the non-ODE baseline models also take the historical context as input, with model-specific encoding choices. See the Appendix for further model implementation details.
To tune our models and compute test error, due to the small dataset size, we use repeated nested cross validation with 3 repeats, 6 outer folds and 4 inner folds. The inner folds are used to tune hyperparameters and outer folds to estimate generalization error; results are averaged over three runs for which the sequences are randomly shuffled. For more details of the evaluation procedure, see the Appendix.
@@ 5.4 Results
In Fig. 4, we show prediction root mean squared error (RMSE) and classification error, calculated using the nested cross validation procedure described above, for each of the considered models as a function of $\alpha$. Here, the classification task is to correctly predict the intervention that yields the highest average glucose among a given set of 3 choices. As we vary $\alpha$, the amount we emphasize the causal loss component of our hybrid loss, we see increasing predictive RMSE for all models; however, the mechanistic and hybrid ODE models are stable across a wide range of $\alpha$ values. On the other hand, for all hybrid and blackbox models, classification error decreases with increasing $\alpha$. Note that the UVA/Padova mechanistic model does not achieve 0 classification error because the mechanistic model does not include all mechanistic components relevant to exercise. For our hybrid ODEs, we see that for $\alpha\in\{0.01,0.1\}$, we achieve very low classification error while maintaining low prediction RMSE. Thus, our hybrid loss provides a win-win for hybrid ODEs: predictive performance exceeding pure mechanistic or blackbox approaches while also having extremely low classification error, again outperforming both the mechanistic and blackbox baselines.
Recall, predicting glucose trajectories during exercise is a very challenging task. Arriving at a highly performant predictive model that also enables valid counterfactual reasoning is thus an important step towards better guiding individuals with T1D to exercise safely.
@@
Figure 4: Predictive loss (RMSE) with its standard error and 10/50/90-th(upper/bar/lower) percentile causal classification error rate for all models, across different values of $\alpha$ (hybrid loss parameter). Within each bar group for a single model (UVA, LP, LPSC, MNODE, BNODE, TCN, LSTM, Tran, S4D), the parameter $\alpha$ increases from left to right as color changes from dark to light, taking the values $\{0,1\mathrm{e}{-4},1\mathrm{e}{-3},1\mathrm{e}{-2},1\mathrm{e}{-1},1\}$ successively. The dashed line in the bottom figure corresponds to the causal classification error of random guessing ($2/3$).
@@ 6 Discussion
[…]
7 Acknowledgements
This work was supported in part by AFOSR Grant FA9550-21-1-0397, ONR Grant N00014-22-1-2110, NSF Grant 2205084, the Stanford Institute for Human-Centered Artificial Intelligence (HAI), the Helmsley Charitable Trust, and the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard. EBF is a Chan Zuckerberg Biohub – San Francisco Investigator. This publication is based on research using data from Jaeb Center for Health Research Foundation that has been made available through Vivli, Inc. Vivli has not contributed to or approved, and is not in any way responsible for, the contents of this publication.
@@ Appendix B Reduced UVA-Padova for DTDSim2
    We removed the whole glucagon system to reduce model variance because most T1D patients do not have exogenous glucagon delivery, and their body’s own glucagon regulation system is often impaired, a common symptom of T1D (Bisgaard Bengtsen and Møller, 2021).
@@ Appendix C MNODE Graph Reduction Heuristic
Appendix D Detailed Data Preparation Procedure
Here we provide a step-by-step procedure of how we pre-processed the T1DEXI
We select patients on open-loop pumps with age under 40 and body mass index (BMI) less than 30. And among the selected patient, we keep exercises that lasted for at least 30 minutes.
For each selected exercise instance, we focus on the time window from 4 hours before the beginning of exercise to 30 minutes after the beginning of exercise. As CGM readings are taken every 5 minutes for all patients in T1DEXI, this 270-minute window should correspond to 54 entries (5 minutes apart) in CGM data. These 54 entries will form the first feature of the time series representation of this exercise and the time stamps of these 54 entries will be the time stamps of the time series. Next we move to extract insulin and meal data. For each time stamp of the time series, we scan the insulin injection/carbohydrate consumption data files and if the patient had injected insulin/consumed carbohydrate between this time stamp and the next, we add the recorded amount to the 2nd/3rd feature of the time series at this time stamp. Note that for insulin we consider both basal insulin and bolus insulin injections. Next, we scan the heart rate/step count file, and set the 4th/5th feature of time series to be the average heart rate/step count recorded between the corresponding time stamp and the next. After this step, one should end up with a time series of 54 time steps and 5 features for each selected exercise instance.
The next major step is to prepare the intervention sets. For each exercise time series, we first create an intervention set containing three copies of the original time series. And then we uniformly randomly (simulated by a numpy default random number generator with seed 2024) apply one of the four following intervention sets to it:

### v2

@@ 4 Related Work
GNS applies the message passing architecture of graph convolution networks (GCN) to physical system simulation (Sanchez-Gonzalez et al., 2018, 2020; Pfaff et al., 2021; Wu et al., 2022; Allen et al., 2023). Poli et al. (2019) connect GNS with blackbox neural ODEs (GNODE) (see also Jin et al., 2022; Bishnoi et al., 2023; Wu et al., 2022; Allen et al., 2023; Li et al., 2022). MNODE focuses on causal relationships between features across time, rather than between spatial interactions; as such, we use directed rather than undirected graphs.
@@ 5.1 Motivation: Safely Managing Exercise in T1D
Type 1 diabetes (T1D) is an autoimmune condition characterized by the destruction of insulin producing beta-cells. To manage glucose concentrations and prevent diabetes-related complications, intensive insulin therapy through externally delivered sources (e.g., insulin pumps) is required. Regular physical activity and exercise lead to numerous health benefits such as increased insulin sensitivity, weight management, and improved psychosocial well-being. However, for individuals with T1D, due to the inability to rapidly decrease circulating insulin concentrations, exercise can also increase the risk of hypoglycemia. Despite technological advancements such as continuous glucose monitoring (CGM) and automated insulin delivery systems, individuals with T1D still face significant challenges with managing glucose concentrations around exercise. Many adults with T1D are not meeting current exercise recommendations of at least 30-minutes of moderate-to-vigorous physical activity per day (Riddell et al., 2017), with fear of exercise-related hypoglycemia a leading barrier (Brazeau et al., 2008). To address these challenges, there is a pressing demand for precise, reliable models that can predict individualized glycemic responses during and after exercise and encourage safe physical activity for all individuals with T1D.
Past efforts on modeling glycemic response to physical activity have focused on developing more intricate mechanistic models (Dalla Man et al., 2009; Liu et al., 2018; Deichmann et al., 2023), rather than devising a performant predictive model. Although glucose prediction using ML and, more recently, deep learning methods has received significant attention, a dearth of papers consider exercise periods (Oviedo et al., 2017) and the few that do do not perform well (Hobbs et al., 2019; Xie & Wang, 2020; Tyler et al., 2022). We aim to leverage hybrid modeling to predict glucose concentrations of an individual with T1D in the first 30 minutes following physical activity, given historical context and expected future covariates. Exercise leads to increased insulin sensitivity post-exercise and this may in turn contribute to a heightened risk of hypoglycemia after activity. The post-exercise period represents both a period of glycemic risk and one in which interventions can be readily applied; our goal is to enable better guidance during this period.
@@ 5.2 Data Preparation
Our data come from the Type 1 Diabetes Exercise Initiative (T1DEXI) (Riddell et al., 2023), which can be requested via https://doi.org/10.25934/PR00008428. This is a real-world study of exercise effects on 497 adults with T1D. Participants in the study were randomly assigned to aerobic, resistance, or interval exercise videos for a total of six sessions over four weeks. Participants were asked to self-report their food intake and exercise habits, while their insulin dosage and relevant physiological responses were recorded by corresponding wearable devices such as insulin pumps, CGM devices, and smart watches.
@@
For selected participants, we filter out exercises shorter than 30 minutes. Then, for each exercise instance, we extract the participant’s metabolic history (basal/bolus insulin delivery, carbohydrate intake, heart rate, step count, and CGM readings) 4 hours before to 30 minutes after the end of exercise to form a 5-dimensional time series. The 4-hour context window was chosen because the effect of most bolus insulin (fast-acting insulin) lasts for around 4 hours. It is common within the T1D community to consider this time range when historical context is relevant in decision making. Finally, we filter out exercise instances with missing heart rate values. We end up with 143 exercise instances from 78 participants. See the Appendix for further data preprocessing details.
Intervention Sets
Our causal loss relies on the introduction of intervention sets. Here, we consider sets of size $K=3$. For our training procedure, for each training post-exercise instance, we create 3 replicates of $(X(\mathcal{T}_{\mathrm{context}})$,$Y(\mathcal{T}_{\mathrm{context}}))$ and append each replicate with an input sequence $X^{(i)}(\mathcal{T}_{\mathrm{pred}})$ from a set selected uniformly at random from the following categories: (1) adding 0/50/100 grams (g) of carbohydrates at the end of exercise; (2) adding 0/2.5/5.0 units of insulin uniformly throughout the first 30 minutes post-exercise; (3) no-change/50g carbs/10.0 units of insulin at the end of exercise; (4) replacing the real post-exercise heart rate trajectory with a prototypical trajectory seen in aerobic/resistance/interval training. All of these modifications are relative to the observed inputs $X(\mathcal{T}_{\mathrm{pred}})$. These intervention sets capture the most integral pieces of well-understood and well-researched domain knowledge, allowing us to compute a reliable true ranking. In particular, increasing levels of carbohydrates (resp., decreasing levels of insulin) progressively increase mean glucose; and increasing exercise intensity (as measured by heart rate) leads to progressively increasing mean glucose during the short exercise period being considered (Aronson et al., 2019; Riddell et al., 2017). For our evaluation, we consider counterfactual simulations defined similarly. For each test exercise instance, we create a set of three instances $(X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}}),X^{(i)}(\mathcal{T}_{\mathrm{pred}}))$ used to generate a counterfactual simulation $\hat{Y}^{(i)}(\mathcal{T}_{\mathrm{pred}})$, $i=1,2,3$. The intervention sets defining $X^{(i)}(\mathcal{T}_{\mathrm{pred}})$ are again drawn uniformly at random. In all cases (train and test), we take score to be average glucose in the 30-minute prediction window.
@@ 5.3 Key Implementation Details
and then computes the gradient via backpropagation. Here, we use subscript instead of parentheses to indicate the transition from the continuous to discrete time. In our implementation, we set $\Delta t$ to be 5 minutes, which is the sampling rate of CGM readings in the T1DEXI data.
@@
Instead of the highly parameterized UVA/Padova S2013 model (Man et al., 2014), our hybrid models build upon a reduced model with the aim of (1) reducing variance and improving generalization performance, and (2) reducing training time. We applied a data-driven reduction method to the full UVA/Padova mechanistic model and its causal graph; see the Appendix. An empirical comparison to hybridizations of the full, non-reduced UVA/Padova model are in the Appendix. The results illustrate that the flexibility of the neural networks in the hybrid models is a satisfactory replacement for many of the full mechanistic model compartments.
@@
We consider 3 hybrid models in the form of Eq. (5) with increasing amounts of flexibility: (1) latent parameter dynamics (LP) defined via $c_{1}=c_{4}=1,c_{2}=c_{3}=0$; (2) latent parameter dynamics plus state closure (LPSC) defined via $c_{1}=c_{2}=c_{4}=1,c_{3}=0$; and (3) mechanistic neural ODE (MNODE) defined via $c_{2}=1,c_{1}=c_{3}=c_{4}=0$. In all of these cases, the adjacency matrices $A_{s}$ and $A_{x}$ are specified based on our reduced UVA/Padova mechanistic model. We implement MNODE by defining a neural network for each output dimension and masking appropriate inputs, though the methods of Chen et al. (2024) could be adapted to our setting to provide a more efficient alternative. We also consider the blackbox neural ODE model (BNODE) of Eq. (3) and tune the latent state dimension. Finally, we provide a set of non-ODE-based blackbox sequence models: the LSTM, Transformer (Trans), temporal convolutional network (TCN), and diagonal approximation to S4 (S4D) (Gu et al., 2022a). For our mechanistic model baseline, we consider the full (not reduced) UVA/Padova S2013 simulator (UVA). We limit all models to have fewer than 25,000 parameters for fair comparison. For all of our ODE-based models, we use a multi-stack LSTM for our initial-condition encoder; the non-ODE baseline models also take the historical context as input, with model-specific encoding choices. See the Appendix for further model implementation details.
To tune our models and compute test error, due to the small dataset size, we use repeated nested cross validation (CV) with 3 repeats, 6 outer folds and 4 inner folds. The inner folds are used to tune hyperparameters and outer folds to estimate generalization error; results are averaged over three runs for which the sequences are randomly shuffled. For more details of the evaluation procedure, see the Appendix.
@@ 5.4 Results
Figure 4: For the synthetic data, predictive loss (RMSE) and standard error (top) and 10/50/90-th(upper/bar/lower) percentile causal classification error rate (bottom). Within each bar group for a single model, the parameter $\alpha$ (hybrid loss parameter) increases from left to right, taking the values $\{0,1\mathrm{e}{-4},1\mathrm{e}{-3},1\mathrm{e}{-2},1\mathrm{e}{-1},1\}$. The dashed line in the bottom figure corresponds to the causal classification error of random guessing ($2/3$).
Synthetic Data
We start by exploring a synthetic data example generated by a simple, single-state ODE with two correlated inputs, mimicking the carb/insulin correlation in the T1DEXI data. In particular, the ODE is specified as:
$dy/dt$ $=-y(t)+x_{1}(t)-x_{2}(t),\quad y(0)=0$
$x_{1}(t)$ $=a\exp(-bt),\;a\sim\mbox{Unif}[1,2],\;b\sim\mbox{Unif}[5,15]$
$x_{2}(t)$ $=1.5x_{1}(t)+\epsilon,\quad\epsilon\sim N(0,1\mathrm{e}{-4}).$
We simulate 600 training, 200 validation, and 200 test sequences and corresponding intervention sets, each selected uniformly at random from: (1) raising $x_{1}$ by 0/+1/+2; (2) raising $x_{2}$ by 0/+1/+2; or (3) no-change/adding +1 to $x_{1}$/adding +1 to $x_{2}$. The sequences are of length 100.
In Fig. 4, we show prediction root mean squared error (RMSE) and classification error, calculated using the nested CV procedure described above, for each of the considered models as a function of $\alpha$. Here, the classification task is to correctly predict the intervention that yields the highest score (taken to be average of $\hat{Y}^{(i)}(\mathcal{T}_{\mathrm{pred}})$) among the set of 3 choices. We see that even when the true mechanistic model is used in the hybrid modeling—and even when that system is very simple—our causal loss is critical in disambiguating the sign of the treatment effects for $x_{1}$ and $x_{2}$. For example, when $\alpha=0$ (fully predictive loss), MNODE (and the blackbox models) have large causal loss because of lack of identifiability between the sign of the treatment effects of $x_{1}$ and $x_{2}$. However, even a small $\alpha>0$ rapidly enables disambiguation between $x_{1}$ and $x_{2}$. These synthetic results illustrate the importance of moving beyond predictive loss, especially when there is significant correlation between inputs; the importance is heightened when working with partial observations, as in our real-data setting below.
T1DEXI Data
We now turn to our results on the real-world T1DEXI data. In Fig. 5, we see that as $\alpha$—the amount we emphasize the causal loss component of our hybrid loss—increases, all hybrid and blackbox models have significant decreases in classification error; however, the hybrid causal error rate decreases more rapidly. Critically, across a broad range of $\alpha$ values, the hybrid models have stable RMSE; the blackbox models appear to be more sensitive to increasing $\alpha$. (The RMSE of all models is impacted for sufficiently large $\alpha$, notably $\alpha=1$ when a purely causal loss is considered.) Note that the UVA/Padova mechanistic model does not achieve zero classification error because the mechanistic model does not include all mechanistic components relevant to exercise. Our hybrid loss provides a win-win for hybrid ODEs: predictive performance exceeding pure mechanistic or blackbox approaches while also having extremely low classification error, again outperforming both the mechanistic and blackbox baselines.
Although our causal loss focuses on intervention ranking, this information alone may bias learning towards causally-grounded models that better capture the direction of treatment effects even outside the domain knowledge encoded in the causal loss. For example, in the Appendix, we define our causal loss in terms of rankings on carbohydrate and insulin intervention sets, but evaluate on insulin-to-carbohydrate ratios, which represents the most common intervention considered by clinicians for T1D. Our experiments demonstrate that even when we train with limited domain knowledge (i.e., the impact of carbohydrates or insulin individually), all models appear to exhibit improved insulin-to-carbohydrate rankings relative to models that do not leverage causal loss.
In the Appendix, we also consider the impact of increasing levels of corrupted domain knowledge where our causal loss is given the incorrect label for the top-ranked intervention, $\mathcal{I}^{*}$, for some fraction of intervention sets. The results indicate robustness to moderate amounts of corruption.
Figure 5: As in Fig. 4, but for T1DEXI data and using a full UVA/Padova mechanistic model baseline and hybridizations of the reduced UVA/Padova simulator for LP, LPSC, and MNODE.
@@ 6 Discussion
We presented H2NCM, a method for integrating domain knowledge not only through a hybrid model, but also through a hybrid loss that encourages causal validity. We consider this in the challenging real-world setting of predicting post-exercise glycemic response in T1D. Our experiments illustrate a win-win where—across a wide range of settings of $\alpha$—our state-of-the-art predictive performance does not drop while the causal validity dramatically improves. In theory, models that can do both tasks well should align better with the true underlying system $M^{\dagger}$ and thus generalize better to unseen data, especially in the presence of distribution shifts, noise, and incompleteness, as our promising initial results in the Appendices suggest.
We assume a known mechanistic model from which to build our hybrid models. One could instead consider softer forms of prior mechanistic knowledge. For example, Wang et al. (2024) specify priors over graphical structures in neural differential equations. Indeed, our hybrid loss may be formulated in a Bayesian context as well, and could aid other modeling frameworks (cf., Takeishi & Kalousis, 2021).
Acknowledgements
This work was supported in part by AFOSR Grant FA9550-21-1-0397, ONR Grant N00014-22-1-2110, NSF Grant 2205084, the Stanford Institute for Human-Centered Artificial Intelligence (HAI), the Helmsley Charitable Trust, and the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard. EBF is a Chan Zuckerberg Biohub – San Francisco Investigator. This publication is based on research using data from Jaeb Center for Health Research Foundation that has been made available through Vivli, Inc. Vivli has not contributed to or approved, and is not in any way responsible for, the contents of this publication. We thank Ke Alexander Wang for helpful discussions on the implementation and training of various models.
Impact Statement
All clinical decision support systems come with some risks, even with a trained human in the loop. If our intervention-ranking system gains the trust of the clinician, they may rely on its recommendations blindly. There are of course cases where it will make errors. In the type 1 diabetes setting, an error of recommending too much insulin may lead to potentially life-threatening hypoglycemia. A benefit of our system, however, is that instead of searching over all possible interventions, we are ranking an expert-provided list of possible interventions, all of which are assumed to be safe alternatives that would not deviate from clinically established standards of care.
@@ Appendix B Reduced UVA-Padova for DTDSim2
    We removed the whole glucagon system to reduce model variance because most T1D patients do not have exogenous glucagon delivery, and their body’s own glucagon regulation system is often impaired, a common symptom of T1D (Bisgaard Bengtsen & Møller, 2021).
@@ Appendix C MNODE Graph Reduction Heuristic
Appendix D Detailed Data Preparation Procedure for T1DEXI
D.1 T1DEXI Exercise Instances
Here we provide a step-by-step procedure of how we pre-processed the T1DEXI dataset described in Section 5.2.
We select patients on open-loop pumps with age under 40 and body mass index (BMI) less than 30. For each selected exercise instance in our dataset, we focus on the time window from 4 hours prior to the end of exercise, to 30 minutes after the end of exercise. We anchor time zero as the end of exercise, so this time window (in minutes) is $[-240,30]$. Since CGM measurements are taken in 5 minute increments, we divide this time window into 5 minute increments, and divide the end time of each interval by 5, so that we obtain 54 discrete time steps: $t=-47,-46,-45,\ldots,4,5,6$, with $t=0$ denoting the end of exercise.
For each exercise instance, we use the following (54-dimensional) features derived from the T1DEXI data: